OpenStudy (anonymous):

Given the function g(x) = −4x + 5, find g(−4)

OpenStudy (anonymous):

(-4 . 4) + 5 = - 16 + 5 = - 11

OpenStudy (jeremy):

(-4.-4) +5 16+5 21

OpenStudy (anonymous):

All you need to is replace the variable x with -4. Substitute x, with -4 and solve.